# Tagged particle fluctuations for TASEP with dynamics restricted by a moving wall

### v1

@@ 5 Functional slow decorrelation
To prove our main result, we will need to apply the slow decorrelation phenomenon, which essentially says that the fluctuations along characteristic lines (macroscopic lines around which the backwards path fluctuates) are the same if one looks at times that differ mesoscopically. Moreover, since the starting formula (2.1) contains the full process, a finite-dimensional slow decorrelation result as in [20, 16] is not enough. We need to upgrade it to a functional slow decorrelation statement as first made in [18], although unlike them, we use tightness directly without a comparison to auxiliary processes.
In our case, we need to argue that what happens in a mesoscopic time at the beginning of the evolution is not creating relevant fluctuations. For this reason, we first show a slow decorrelation result for the following rescaled tagged particle processes:
@@
=\ = $\Bigl{(}1-2\sqrt{\tfrac{\alpha T}{t}}\Bigr{)}t-\Bigl{(}1-\sqrt{\tfrac{\alpha T}{t}}-\tfrac{\alpha}{\alpha_{i}}\sqrt{\tfrac{t}{\alpha T}}\Bigr{)}T^{\nu}+{\cal O}(T^{2\nu-5/3})$
=\ = $\Bigl{(}1-2\sqrt{\tfrac{\alpha T}{t}}\Bigr{)}t-\Bigl{(}1-2\sqrt{\tfrac{\alpha}{\alpha_{i}}}\Bigr{)}T^{\nu}+{\cal O}(T^{\nu-2/3})$
with $T^{\nu-2/3}=o(T^{1/3})$. Thus, pointwise slow decorrelation can be obtained by the usual method proposed in [16], see also [12, 21, 24]. By (A.4), for each fixed $\tau$ (recall that then $t={\cal O}(T)$) we have
$\tfrac{x_{\alpha T}(t)-\Bigl{(}1-2\sqrt{\tfrac{\alpha T}{t}}\Bigr{)}t}{-\tilde{c}_{1}^{i}T^{1/3}}\Rightarrow$ $F_{2},$ (5.6)
$\tfrac{y^{x_{i}}_{\alpha T-{\alpha\over\alpha_{i}}T^{\nu}}(T^{\nu},t)-\biggl{(}1-2\sqrt{\tfrac{\alpha T-{\alpha\over\alpha_{i}}T^{\nu}}{t-T^{\nu}}}\biggr{)}(t-T^{\nu})}{-\tilde{c}_{1}^{i}T^{1/3}}\Rightarrow$ $F_{2},$ (5.7)
$\tfrac{x_{{\alpha\over\alpha_{i}}T^{\nu}}(T^{\nu})-\bigl{(}1-2\sqrt{\tfrac{\alpha}{\alpha_{i}}}\bigr{)}T^{\nu}}{-\tilde{c}_{1}^{i}T^{\nu/3}}\Rightarrow$ $F_{2}$ (5.8)
@@
The second summand on the right hand side converges to 00 in probability since we scale by $T^{1/3}$ instead of $T^{\nu/3}$. Hence, Slutzky’s theorem, see Theorem 13.18 of [38], implies that both sides of the inequality converge weakly to the same limit. By Lemma 4.1 of [5], their difference converges to 00 in probability. This yields the pointwise slow decorrelation result: let $t=\alpha_{i}T-\tilde{c}_{2}^{i}\tau T^{2/3}$ with $\tau\in[-\varkappa,\varkappa]$ fixed. Then, for each $\varepsilon>0$
@@
From (5.10), we obtain convergence of $(\tilde{X}^{i}_{T}(\tau)-\tilde{X}^{\nu,i}_{T}(\tau))$ to 00 in the sense of finite-dimensional distributions. By Lemma 2.2, $(\tilde{X}_{T}^{i}(\tau))$ is tight. Since $x_{i}$ as well as $T^{\nu}$ do not depend on $\tau$, we have
@@
$(\tilde{X}^{\nu,i}_{T}(\tau))\overset{(d)}{=}\biggl{(}\tfrac{x^{\text{step}}_{\alpha\tilde{T}}(\alpha_{i}\tilde{T}-\tilde{c}_{2}^{i}c_{\tilde{T}}\tau\tilde{T}^{2/3})-\tilde{\mu}^{i}(c_{\tilde{T}}\tau,\tilde{T})}{-\tilde{c}_{1}^{i}\tilde{T}^{1/3}}\ d_{\tilde{T}}\biggr{)}\overset{(d)}{=}(\tilde{X}_{\tilde{T}}^{i}(c_{\tilde{T}}\tau)d_{\tilde{T}})$ (5.12)
with $c_{\tilde{T}}=T^{2/3}\tilde{T}^{-2/3}\to 1$ and $d_{\tilde{T}}=\tilde{T}^{1/3}T^{-1/3}\to 1$ as $T\to\infty$. In [8], it is shown that $(\tilde{X}_{T}^{i}(\tau))$ satisfies the conditions of Theorem 15.5 of [6]. Then, tightness of the process is preserved under the modifications of multiplying $\tau$ as well as the process itself by converging, deterministic sequences $c_{\tilde{T}},d_{\tilde{T}}$. This yields tightness of $(\tilde{X}^{\nu,i}_{T}(\tau))$. If two processes fulfil the conditions of Theorem 15.5 of [6], then the same holds true for their difference. Thus, $(\tilde{X}^{i}_{T}(\tau)-\tilde{X}^{\nu,i}_{T}(\tau))$ is tight and converges to 00 in the sense of finite-dimensional distributions. By Theorem 15.1 of [6], this gives weak convergence of $(\tilde{X}^{i}_{T}(\tau)-\tilde{X}^{\nu,i}_{T}(\tau))$ to 00 in the space of càdlàg functions on $[-\varkappa,\varkappa]$. As the limit is continuous, we also have weak convergence with respect to the uniform topology. This concludes the proof of Proposition 5.1. ∎
@@ 6 Proof of the main results
and to extend Proposition 4.1 from fixed times to the time intervals $[t_{l}^{i},t_{r}^{i}]$. For this, we define regions [^4] For notational simplicity, we do not restrict the time intervals as in (4.1) because we consider particles at different times here. However, it is still enough to construct the processes up to a finite time (like $T$), so we could replace $\infty$ by it in the definition of the regions.
@@
and TASEPs $\tilde{x}^{r,i}(t),\tilde{x}^{l,i}(t)$ analogously as in Proposition 4.1: the processes $x^{\text{step}}(t),\tilde{x}^{r,i}(t),\tilde{x}^{l,i}(t)$ have the same (step) initial condition and share their jump attempts, except that $\tilde{x}^{r,i}(t)$ has density 00 in $({\cal C}_{i}^{r})^{c}$ and $\tilde{x}^{l,i}(t)$ has density $1$ in $({\cal C}_{i}^{l})^{c}$. We set $N_{i}=\alpha T-\tfrac{\alpha}{\alpha_{i}}T^{\nu}$.
@@
[…]
@@
Then, $x^{\text{step}}_{N}(t_{l}-T^{\nu})>x^{\text{step},h}_{M_{l}}(t_{l}-T^{\nu})$ is equivalent to $x^{\text{step}}_{N}(t-T^{\nu})>x^{\text{step},h}_{M_{l}}(t-T^{\nu})$ for all $t\in[t_{l},t_{r}]$ since once a particle jumped over a hole, their paths will not cross again. As in the proof of Proposition 4.1, Lemma A.1 yields
@@
[…]
for some constant $\tilde{c}>0$ and $T$ large enough. As stated in the proof of Proposition 4.1, the last probability can be bounded by $Ce^{-cK}$ by Proposition 4.2, for $K$ large enough. Thus, $\mathbb{P}(E_{l}^{c})\leq Ce^{-cK}$. ∎
@@
Figure 7: This image shows (6.13) for $t=\alpha_{0}T$ and $t=\alpha_{1}T$, under the assumption that we localized the backwards paths starting at $x^{\text{step},x_{i}}_{N_{i}}(T^{\nu},t)$ completely. The evolution of $x_{\alpha T}(t)$ corresponds to the black line and those of $x^{\text{step},x_{i}}_{N_{i}}(T^{\nu},t)$ correspond to the green respectively blue lines. Their backwards paths starting at times $\alpha_{i}T$, depicted by the orange respectively red lines, are contained within ${\cal O}(T^{2/3})-$neighbourhoods of $\Bigl{(}1-2\sqrt{\tfrac{\alpha}{\alpha_{i}}}\Bigr{)}t$ and do not meet with probability converging to $1$ because the distance of $x_{0}$ and $x_{1}$ is of order $T^{\nu}\gg T^{2/3}$.
@@
where $K_{0}\leq K={\cal O}(T^{\varepsilon})$ for some constant $K_{0}=K_{0}(\varkappa)>0$. The constants $K_{0},C,c>0$ can be chosen uniformly for all times $T$ large enough. The process $\tilde{x}^{r,i}(T^{\nu},t)$ has density 00 in $({\cal D}_{i}^{r})^{c}$ and the process $\tilde{x}^{l,i}(T^{\nu},t)$ has density $1$ in $({\cal D}_{i}^{l})^{c}$.
@@
jointly, and $\tilde{x}^{r,i}(T^{\nu},t)$ has density 00 in the region
@@
The existence of an event $E_{i}$ with the desired properties follows from Proposition 6.2. For
@@
Thus, we can replace $E_{i}$ by $E_{i}\cap B_{i}$ and, given this event, obtain
@@
=\ = $\mathbb{P}((\tilde{x}^{r,0}_{N_{0}}(T^{\nu},t))_{t\in[t_{l}^{0},t_{r}^{0}]}\in A_{0},(\tilde{x}^{l,1}_{N_{1}}(T^{\nu},t))_{t\in[t_{l}^{1},t_{r}^{1}]}\in A_{1}|E)$
=\ = $\mathbb{P}((x^{\textup{step},x_{0}}_{N_{0}}(T^{\nu},t))_{t\in[t_{l}^{0},t_{r}^{0}]}\in A_{0}|E)\ \mathbb{P}((x^{\textup{step},x_{1}}_{N_{1}}(T^{\nu},t))_{t\in[t_{l}^{1},t_{r}^{1}]}\in A_{1}|E).$
@@
$A^{T}_{\varepsilon}\coloneqq\{|\hat{X}_{T}^{\nu,i}(\tau)-\tilde{X}_{T}^{i}(\tau)|\leq\varepsilon\ \forall\tau\in[-\varkappa,\varkappa],i\in\{0,1\}\},$ (6.24)
@@
[…]
Let $[-\varkappa,\varkappa]$ be an interval with $\varkappa\geq 1$ fixed as $T\to\infty$. Then, for $T$ large enough it holds
$[-\varkappa,\varkappa]\subseteq[-\varepsilon(\tilde{c}_{2}^{0})^{-1}T^{1/3},\varepsilon(\tilde{c}_{2}^{0})^{-1}T^{1/3}]\cap[-\varepsilon(\tilde{c}_{2}^{1})^{-1}T^{1/3},\varepsilon(\tilde{c}_{2}^{1})^{-1}T^{1/3}].$ (6.32)
[…]
@@
for $\varkappa$ large enough, and the right hand side converges to 00 as $\varkappa\to\infty$. This yields (6.36). For representations of the probabilities as Fredholm determinants, refer to Theorem 1.19 and Equation (3.4) of [46] and Theorem 2 of [19].
@@
$J\coloneqq[0,T]\setminus([\alpha_{0}T-\tilde{c}_{2}^{0}\varkappa T^{2/3},\alpha_{0}T+\tilde{c}_{2}^{0}\varkappa T^{2/3}]\cup[\alpha_{1}T-\tilde{c}_{2}^{1}\varkappa T^{2/3},\alpha_{1}T+\tilde{c}_{2}^{1}\varkappa T^{2/3}])$ (6.38)
@@
$E_{T,\varkappa}^{i}\coloneqq\{\tilde{X}_{T}^{i}(\tau)+\tau^{2}-g_{T}^{i}(\tau)\leq S(\tilde{c}_{1}^{i})^{-1}\forall|\tau|\leq\varkappa\},\ i\in\{0,1\},$ (6.39)
$A_{T,\varkappa}\coloneqq\{x_{\alpha T}(t)\geq\xi T-ST^{1/3}-f(T-t)\forall t\in J\}.$
@@
[…]
In Lemma 6.5, we establish:
@@
Finally, we formulate the two auxiliary lemmata needed in the previous proof.
@@
As this convergence is obtained analogously in the case of one wall influence in [8], we do not repeat the details here. The idea is to split the interval $J$ up into further subintervals. For the intervals $[0,\alpha T]$ and $[\alpha T,(\alpha+\delta)T]$ with some small $\delta>0$, one utilizes Assumption 2.3 (a) and the fact that it holds $x_{\alpha T}(t)\geq-\alpha T$ for all times $t\geq 0$. For the remaining times outside of ${\cal O}(T^{2/3+\sigma})-$neighbourhoods of $\alpha_{0}T$ and $\alpha_{1}T$, where $\sigma\in(0,\tfrac{1}{6})$, one can argue by Assumption 2.3 (a) and (b) as well as Lemma A.1. Finally, inside of the ${\cal O}(T^{2/3+\sigma})-$neighbourhoods, one utilizes Assumption 2.3 (b) and the comparison to stationary TASEP by Proposition 3.15. We refer to Lemma 4.11, Lemma 4.12, Lemma 4.13 and Lemma 4.14 of [8].
Lemma 6.6.
[…]
$\sup_{\tau\in I}\{\mathcal{A}_{2}(\tau)-g(\tau)\}$ (6.45)
is continuous.
Proof.
[…]
and the latter converges to $1$ as $M\to\infty$.

### v2

@@ 5 Functional slow decorrelation
The fluctuations of particle positions at a macroscopic time $t$ are of order $t^{1/3}$. However, fluctuations along characteristic lines (macroscopic lines around which the backwards path fluctuates), at times differing on a mesoscopic scale, are the same up to $o(t^{1/3})$. This is called slow decorrelation phenomenon and will be needed in the proof of our main result. Since the starting formula (2.1) contains the full process, a finite-dimensional slow decorrelation result as in [20, 16] is not enough. We need to upgrade it to a functional slow decorrelation statement as first made in [18], although unlike them, we use tightness directly without a comparison to auxiliary processes.
In our case, we need to argue that what happens in a mesoscopic time at the beginning of the evolution is not creating relevant fluctuations as they are in $o(T^{1/3})$. For this reason, we first show a slow decorrelation result for the following rescaled tagged particle processes:
@@
$=$ $\Bigl{(}1-2\sqrt{\tfrac{\alpha T}{t}}\Bigr{)}t-\Bigl{(}1-\sqrt{\tfrac{\alpha T}{t}}-\tfrac{\alpha}{\alpha_{i}}\sqrt{\tfrac{t}{\alpha T}}\Bigr{)}T^{\nu}+{\cal O}(T^{2\nu-5/3})$
$=$ $\Bigl{(}1-2\sqrt{\tfrac{\alpha T}{t}}\Bigr{)}t-\Bigl{(}1-2\sqrt{\tfrac{\alpha}{\alpha_{i}}}\Bigr{)}T^{\nu}+{\cal O}(T^{\nu-2/3})$
with $T^{\nu-2/3}=o(T^{1/3})$. Thus, pointwise slow decorrelation (see (5.10) below) can be obtained by the usual method proposed in [16], see also [12, 21, 24]. By (A.4), for each fixed $\tau$ (recall that then $t={\cal O}(T)$) we have
$\tfrac{x_{\alpha T}(t)-\Bigl{(}1-2\sqrt{\tfrac{\alpha T}{t}}\Bigr{)}t}{-\tilde{c}_{1}^{i}T^{1/3}}\Rightarrow\$$F_{2},$ (5.6)
$\tfrac{y^{x_{i}}_{\alpha T-{\alpha\over\alpha_{i}}T^{\nu}}(T^{\nu},t)-\biggl{(}1-2\sqrt{\tfrac{\alpha T-{\alpha\over\alpha_{i}}T^{\nu}}{t-T^{\nu}}}\biggr{)}(t-T^{\nu})}{-\tilde{c}_{1}^{i}T^{1/3}}\Rightarrow\$$F_{2},$ (5.7)
$\tfrac{x_{{\alpha\over\alpha_{i}}T^{\nu}}(T^{\nu})-\bigl{(}1-2\sqrt{\tfrac{\alpha}{\alpha_{i}}}\bigr{)}T^{\nu}}{-\tilde{c}_{1}^{i}T^{\nu/3}}\Rightarrow\$$F_{2}$ (5.8)
@@
The second summand on the right hand side converges to 0 in probability since we scale by $T^{1/3}$ instead of $T^{\nu/3}$. Hence, Slutzky’s theorem, see Theorem 13.18 of [38], implies that both sides of the inequality converge weakly to the same limit. By Lemma 4.1 of [5], their difference converges to 0 in probability. Explicitly, this yields the pointwise slow decorrelation result: let $t=\alpha_{i}T-\tilde{c}_{2}^{i}\tau T^{2/3}$ with $\tau\in[-\varkappa,\varkappa]$ fixed. Then, for each $\varepsilon>0$
@@
From (5.10), we obtain convergence of $(\tilde{X}^{i}_{T}(\tau)-\tilde{X}^{\nu,i}_{T}(\tau))$ to 0 in the sense of finite-dimensional distributions. By Lemma 2.2, $(\tilde{X}_{T}^{i}(\tau))$ is tight. Since $x_{i}$ as well as $T^{\nu}$ do not depend on $\tau$, we have
@@
[…]
with $c_{\tilde{T}}=T^{2/3}\tilde{T}^{-2/3}\to 1$ and $d_{\tilde{T}}=\tilde{T}^{1/3}T^{-1/3}\to 1$ as $T\to\infty$. In [8], it is shown that $(\tilde{X}_{T}^{i}(\tau))$ satisfies the conditions of Theorem 15.5 of [6]. Then, tightness of the process is preserved under the modifications of multiplying $\tau$ as well as the process itself by converging, deterministic sequences $c_{\tilde{T}},d_{\tilde{T}}$. This yields tightness of $(\tilde{X}^{\nu,i}_{T}(\tau))$. If two processes fulfil the conditions of Theorem 15.5 of [6], then the same holds true for their difference. Thus, $(\tilde{X}^{i}_{T}(\tau)-\tilde{X}^{\nu,i}_{T}(\tau))$ is tight and converges to 0 in the sense of finite-dimensional distributions. By Theorem 15.1 of [6], this gives weak convergence of $(\tilde{X}^{i}_{T}(\tau)-\tilde{X}^{\nu,i}_{T}(\tau))$ to 0 in the space of càdlàg functions on $\varkappa,\varkappa]$. As the limit is continuous, we also have weak convergence with respect to the uniform topology. This concludes the proof of Proposition 5.1. ∎
@@ 6 Proof of the main results
and to extend Proposition 4.1 from fixed times to the time intervals $t_{l}^{i},t_{r}^{i}]$. For this, we define regions [^7] For notational simplicity, we do not restrict the time intervals as in (4.1) because we consider particles at different times here. However, it is still enough to construct the processes up to a finite time (like $T$), so we could replace $\infty$ by it in the definition of the regions.
@@
and TASEPs $\tilde{x}^{r,i}(t),\tilde{x}^{l,i}(t)$ analogously as in Proposition 4.1: the processes $x^{\text{step}}(t),\tilde{x}^{r,i}(t),\tilde{x}^{l,i}(t)$ have the same (step) initial condition and share their jump attempts, except that $\tilde{x}^{r,i}(t)$ has density 0 in $({\cal C}_{i}^{r})^{c}$ and $\tilde{x}^{l,i}(t)$ has density $1$ in $({\cal C}_{i}^{l})^{c}$. We set $N_{i}=\alpha T-\tfrac{\alpha}{\alpha_{i}}T^{\nu}$.
@@
$\leq$ $\mathbb{P}\Bigl{(}\exists s\in[0,t_{r}-T^{\nu}]:\ x^{\text{step}}_{N(t_{r}-T^{\nu}\downarrow s)}(s)\geq\Bigl{(}1-2\sqrt{\tfrac{N}{t_{r}-T^{\nu}}}\Bigr{)}s+(K-\tilde{c}\varkappa)T^{2/3}\Bigr{)}$
@@
Then, $x^{\text{step}}_{N}(t_{l}-T^{\nu})>x^{\text{step},h}_{M_{l}}(t_{l}-T^{\nu})$ is equivalent to $x^{\text{step}}_{N}(t-T^{\nu})>x^{\text{step},h}_{M_{l}}(t-T^{\nu})$ for all $t\in[t_{l},t_{r}]$ since once a particle jumped over a hole, their paths will not cross again. As in the proof of Proposition 4.1 (see (4.80) and use that $t_{l}-T^{\nu}$ is a macroscopic time), Lemma A.1 yields
@@
$\leq$ $\mathbb{P}\Bigl{(}\exists s\in[0,t_{r}-T^{\nu}]:\ x^{\text{step},h}_{M_{l}(t_{r}-T^{\nu}\downarrow s)}(s)\leq\Bigl{(}1-2\sqrt{\tfrac{\alpha}{\alpha_{i}}}\Bigr{)}s-KT^{2/3}\Bigr{)}$
$\leq$ $\mathbb{P}\Bigl{(}\exists s\in[0,t_{r}-T^{\nu}]:\ x^{\text{step},h}_{M_{l}(t_{r}-T^{\nu}\downarrow s)}(s)+\Bigl{(}1-2\sqrt{\tfrac{M_{l}}{t_{r}-T^{\nu}}}\Bigr{)}s\leq-(K-\tilde{c}\varkappa)T^{2/3}\Bigr{)}$
for some constant $\tilde{c}>0$ and $T$ large enough. As stated in the proof of Proposition 4.1 (see (4.81)), the last probability can be bounded by $Ce^{-cK}$ by Proposition 4.2, for $K$ large enough. Thus, $\mathbb{P}(E_{l}^{c})\leq Ce^{-cK}$. ∎
@@
Figure 7: This image shows (6.13) for $t=\alpha_{0}T$ and $t=\alpha_{1}T$, under the assumption that we localized the backwards paths starting at $x^{\text{step},x_{i}}_{N_{i}}(T^{\nu},t)$ completely. The evolution of $x_{\alpha T}(t)$ corresponds to the black line and those of $x^{\text{step},x_{i}}_{N_{i}}(T^{\nu},t)$ correspond to the green respectively blue lines. Their backwards paths starting at times $\alpha_{i}T$, depicted by the orange respectively red lines, are contained within ${\cal O}(T^{2/3})$-neighbourhoods of $\Bigl{(}1-2\sqrt{\tfrac{\alpha}{\alpha_{i}}}\Bigr{)}t$ and do not meet with probability converging to 11 because the distance of $x_{0}$ and $x_{1}$ is of order $T^{\nu}\gg T^{2/3}$.
@@
where $K_{0}\leq K={\cal O}(T^{\varepsilon})$ for some constant $K_{0}=K_{0}(\varkappa)>0$. The constants $K_{0},C,c>0$ can be chosen uniformly for all times $T$ large enough. The process $\tilde{x}^{r,i}(T^{\nu},t)$ has density 0 in $({\cal D}_{i}^{r})^{c}$ and the process $\tilde{x}^{l,i}(T^{\nu},t)$ has density $1$ in $({\cal D}_{i}^{l})^{c}$.
@@
jointly, and $\tilde{x}^{r,i}(T^{\nu},t)$ has density 0 in the region
@@
The existence of an event with the desired properties follows from Proposition 6.2. We denote the event $E_{i}$ from Proposition 6.2 by $\tilde{E}_{i}$ here. For
@@
Thus, we define $E_{i}=\tilde{E}_{i}\cap B_{i}$ and, given this event, obtain
@@
$=$ $\mathbb{P}((\tilde{x}^{r,0}_{N_{0}}(T^{\nu},t))_{t\in[t_{l}^{0},t_{r}^{0}]}\in A_{0},(\tilde{x}^{l,1}_{N_{1}}(T^{\nu},t))_{t\in[t_{l}^{1},t_{r}^{1}]}\in A_{1}|E)$
$=$ $\mathbb{P}((x^{\textup{step},x_{0}}_{N_{0}}(T^{\nu},t))_{t\in[t_{l}^{0},t_{r}^{0}]}\in A_{0}|E)\ \mathbb{P}((x^{\textup{step},x_{1}}_{N_{1}}(T^{\nu},t))_{t\in[t_{l}^{1},t_{r}^{1}]}\in A_{1}|E).$
Remark 6.5.
@@
The proofs below require the following result.
Lemma 6.6.
Let $\mathcal{A}_{2}$ denote an Airy2 process and let $I\subseteq\mathbb{R}$ be a bounded interval. If $g$ is a càdlàg and piecewise continuous function satisfying $g(\tau)\geq-M+\tfrac{\tau^{2}}{2},\tau\in\mathbb{R},$ for some constant $M\in\mathbb{R}$, then the random variable
$\sup_{\tau\in I}\{\mathcal{A}_{2}(\tau)-g(\tau)\}$ (6.23)
is continuous.
Proof.
For $g$ being continuous, this result has been proven in [18], refer to the equations (103), (104) and Remark 2. There, one can pass from compact intervals $I$ to compactly contained ones since $g$ is continuous and the Airy2 process has continuous sample paths. We want to point out that $\sup_{\tau\in\overline{I}}\{\mathcal{A}_{2}(\tau)-g(\tau)\}$ is finite almost surely as the same already holds for $\sup_{\tau\in\overline{I}}\{\mathcal{A}_{2}(\tau)\}$: for each $M>0$ and $n\in\mathbb{N}$, we have
and the latter converges to $1$ as $M\to\infty$.
The first difference converges to zero by the induction hypothesis, and the second one by the result for continuous functions.
@@
$A^{T}_{\varepsilon}:=\{|\hat{X}_{T}^{\nu,i}(\tau)-\tilde{X}_{T}^{i}(\tau)|\leq\varepsilon\ \forall\tau\in[-\varkappa,\varkappa],i\in\{0,1\}\},$ (6.27)
@@
$\Leftrightarrow$ $\tilde{X}_{T}^{i}(\tau)+\tau^{2}-g_{T}^{i}(\tau)\leq S(\tilde{c}_{1}^{i})^{-1}\textrm{ for all }|\tau|\leq\varepsilon(\tilde{c}_{2}^{i})^{-1}T^{1/3}.$
Let $\varkappa,\varkappa]$ be an interval with $\varkappa\geq 1$ fixed as $T\to\infty$. Then, for $T$ large enough it holds
$\varkappa,\varkappa]\subseteq[-\varepsilon(\tilde{c}_{2}^{0})^{-1}T^{1/3},\varepsilon(\tilde{c}_{2}^{0})^{-1}T^{1/3}]\cap[-\varepsilon(\tilde{c}_{2}^{1})^{-1}T^{1/3},\varepsilon(\tilde{c}_{2}^{1})^{-1}T^{1/3}].$ (6.35)
By Assumption 2.3 (b), the sequences $(g_{T}^{0})$ and $(g_{T}^{1})$ converge uniformly on $\varkappa,\varkappa]$ to $g_{0}$ and $g_{1}$ respectively. Moreover, we obtain weak convergence of $(\tilde{X}_{T}^{i}(\tau)+\tau^{2})$ to $(\mathcal{A}_{2}^{i}(\tau))$ in the space of càdlàg functions on compact intervals from Lemma 2.2, where $\mathcal{A}_{2}^{i}$ denotes an Airy2 process. Therefore, on the interval $\varkappa,\varkappa]$, Theorem 4.4 of [6] yields $(\tilde{X}_{T}^{i}(\tau)+\tau^{2},g_{T}^{i}(\tau))\Rightarrow(\mathcal{A}_{2}^{i}(\tau),g_{i}(\tau))$. In general, the pointwise addition of functions in $\mathbb{D}([-\varkappa,\varkappa])$ is not a continuous mapping [6, p. 123, Problem 3]. Still, since the limits are independent and the Airy2 process has continuous sample paths, addition indeed preserves convergence in this case [50, Section 4]. Therefore, by the continuous mapping theorem, see Theorem 5.1 of [6], it holds
@@
for $\varkappa$ large enough, and the right hand side converges to 0 as $\varkappa\to\infty$. This yields (6.39). For representations of the probabilities as Fredholm determinants, refer to Theorem 1.19 and Equation (3.4) of [46] and Theorem 2 of [19].
@@
$J:=[0,T]\setminus([\alpha_{0}T-\tilde{c}_{2}^{0}\varkappa T^{2/3},\alpha_{0}T+\tilde{c}_{2}^{0}\varkappa T^{2/3}]\cup[\alpha_{1}T-\tilde{c}_{2}^{1}\varkappa T^{2/3},\alpha_{1}T+\tilde{c}_{2}^{1}\varkappa T^{2/3}])$ (6.41)
@@
$E_{T,\varkappa}^{i}=\{\tilde{X}_{T}^{i}(\tau)+\tau^{2}-g_{T}^{i}(\tau)\leq S(\tilde{c}_{1}^{i})^{-1}\forall|\tau|\leq\varkappa\},\ i\in\{0,1\},$ (6.42)
$A_{T,\varkappa}=\{x_{\alpha T}(t)\geq\xi T-ST^{1/3}-f(T-t)\forall t\in J\}.$
@@
As the limit distribution above is a product measure, we can choose the Airy2 processes $\mathcal{A}_{2}^{0},\mathcal{A}_{2}^{1}$ to be independent, such that the joint distribution of $\sup_{\tau\in\mathbb{R}}\{\mathcal{A}_{2}^{0}(\tau)-g_{0}(\tau)\}$ and $\sup_{\tau\in\mathbb{R}}\{\mathcal{A}_{2}^{1}(\tau)-g_{1}(\tau)\}$ takes this form. This is of no direct significance for our result, but it is meant to emphasize the asymptotic independence required to obtain the limit distribution.
In Lemma 6.7 below the proof, we establish:
@@
Finally, we formulate the auxiliary lemma needed in the previous proof.
@@
As this convergence is obtained analogously in the case of one wall influence in [8], we do not repeat the details here.
The idea is to split the interval $J$ up into further subintervals. For the intervals $0,\alpha T]$ and $\alpha T,(\alpha+\delta)T]$ with some small $\delta>0$, one utilizes Assumption 2.3 (a) and the fact that it holds $x_{\alpha T}(t)\geq-\alpha T$ for all times $t\geq 0$. For the remaining times outside of ${\cal O}(T^{2/3+\sigma})$-neighbourhoods of $\alpha_{0}T$ and $\alpha_{1}T$, where $\sigma\in(0,\tfrac{1}{6})$, one can argue by Assumption 2.3 (a) and (b) as well as Lemma A.1. Finally, inside of the ${\cal O}(T^{2/3+\sigma})$-neighbourhoods, one utilizes Assumption 2.3 (b) and the comparison to stationary TASEP by Proposition 3.15. We refer to Lemma 4.11, Lemma 4.12, Lemma 4.13 and Lemma 4.14 of [8].